SOHO DAILY REPORT Operations Day -- 343 1995 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-951210 =========================================================== DATA LOSSES: To be updated this week PASS ACTIVITIES SUMMARY: SVM Activities: * FOT continued adjust the OBT. The Fine phase correction continued. OBT adjustment was completed with a near zero time delta. The time will continue to be monitored. * Min/Max monitoring tables were dumped and reset * A new guide star was acquired. The first attempt failed. A second try with a larger search window located new stars. * Monitored LASCO QTLI1A, QTLI2A. Temps appeared low. After consulting with the PI team, the T3 limits on the spacecraft have been lowered to 0C to avoid a thermal reconfiguration. * Monitored SUMER QTSI1A. Temp appeared high. After consulting with the PI team, the T4 limit on the spacecraft has been raised to 35C to avoid a thermal reconfiguration. * OCD 46, 64, 66, 68, 69 completed. PLM Activities: * VIRGO delayed command load uplinked * SUMER switched on. Later, an OCD was implemented to switch on a substitution heater along with the experiment to maintain thermal conditions. * CEPAC Commissioning Block #2 performed. After review, the CEPAC activities were completed later in the pass. * GOLF turned on. NRT was provided for GOLF initial procedures. * SWAN NRT was used for adjustment to the instrument and troubleshooting of a motor anomaly that occurred previously * UVCS NRT was used for adjustment to the instrument. Starting tomorrow, a more detailed and standardized report will be available from the FOT. --------- end of report ----------